# RocketKV: Accelerating Long-Context LLM Inference via Two-Stage KV Cache Compression

### v1

@@ Abstract
Transformer-based Large Language Models rely critically on KV cache to efficiently handle extended contexts during the decode phase. Yet, the size of the KV cache grows proportionally with the input length, burdening both memory bandwidth and capacity as decoding progresses. To address this challenge, we present RocketKV, a training-free KV cache compression strategy designed specifically to reduce both memory bandwidth and capacity demand of KV cache during the decode phase. RocketKV contains two consecutive stages. In the first stage, it performs coarse-grain KV cache eviction on the input sequence tokens with SnapKV++, a method improved upon SnapKV by introducing adaptive pooling size and full compatibility with grouped-query attention. In the second stage, it adopts a hybrid attention method to conduct fine-grain top-k sparse attention, approximating the attention scores by leveraging both head and sequence dimensional reductions. Combining these two stages, RocketKV achieves significant KV cache fetching bandwidth and storage savings while maintaining comparable accuracy to full KV cache attention. We show that RocketKV provides end-to-end speedup by up to 3$\times$ as well as peak memory reduction by up to 31% in the decode phase on an NVIDIA H100 GPU compared to the full KV cache baseline, while achieving negligible accuracy loss on a variety of long-context tasks.
Payman Behnam* $\circ$ 12 Yaosheng Fu*1 Ritchie Zhao1 Po-An Tsai1 Zhiding Yu1 Alexey Tumanov2
footnotetext: Corresponding authors: Payman Behnam, payman.behnam@gatech.edu.edu; Yaosheng Fu, yfu@nvidia.com
*: Equal contribution
$\circ$: Work done during Payman’s internship at NVIDIA Research
1NVIDIA, Santa Clara, USA
2Georgia Institute of Technology, Atlanta, USA
@@ 1 Introduction
Long-context Large Language Models (LLMs) have enhanced the capabilities of LLMs across diverse tasks that require processing extensive amounts of text. Examples of such applications are document summarization, legal document analysis, coding assistants, conversational AI, etc [24, 22, 12].
@@
In transformer-based LLM [33] inference, the key-value cache (KV cache)—which stores past attention keys and values to avoid recomputation—becomes a major bottleneck during the decode phase, as its size scales linearly with both the sequence length and batch size. For example, the Llama3.1-70B-Instruct [1] model with a batch size of 32, and a context length of 32K requires around 320GB of KV cache storage at FP16 precision, which even advanced hardware [23, 15, 3] can hardly handle.
@@
Fortunately, previous work [37, 31, 25, 20, 28, 9] has shown that only a small subset of KV tokens is required at each decode step to maintain accuracy. Therefore, if those KV tokens can be accurately predicted in advance, dense attention operations can be replaced with sparse attention operations with significant memory bandwidth and capacity improvement. These methods often fall into two categories: 1) permanent KV token eviction, and 2) dynamic KV token selection. The former results in both memory bandwidth and storage savings, but could lead to noticeable accuracy loss if KV tokens dropped earlier are needed by later decode steps. The latter avoids this shortcoming by keeping all KV tokens in the memory and dynamically selecting a subset each time. Hence, it only results in memory bandwidth savings, but often requires extra memory storage overhead for auxiliary data.
@@
Aiming to achieve better trade-offs among memory bandwidth, capacity, and model accuracy, we propose RocketKV, a two-stage KV cache compression method that combines permanent KV token eviction with dynamic KV token selection to accelerate the decode phase of LLM inference. In the first stage, we propose SnapKV++ for coarse-grain permanent KV token eviction on the input prompt. SnapKV++ improves upon SnapKV [20] with adaptive pooling size and full compatibility with grouped-query attention (GQA). In the second stage, RocketKV performs fine-grain dynamic KV token selection with a hybrid attention mechanism which estimates KV token indices with top-k attention scores by leveraging both head and sequence dimensional reductions. Combining these two stages together, RocketKV achieves significant memory bandwidth and storage savings with negligible accuracy loss across a wide variety of models and downstream tasks.
In summary, the contributions of the paper can be summarized as follows:
@@
  • We analyze and identify the limitations of existing KV token dropping methods and then propose RocketKV with two-stage KV cache compression.
  • In the first stage, we propose SnapKV++ for prompt KV cache eviction that improves upon SnapKV by introducing adaptive pooling size and full compatibility with GQA. In the second stage, we propose hybrid attention for dynamic KV token selection at each decode step.
@@
  • We conduct a comprehensive evaluation of RocketKV on a wide variety of models and downstream tasks. RocketKV consistently demonstrates comparable accuracy with full-KV attention while maintaining KV token budgets as low as 256.
  • We evaluate the efficiency of RocketKV on an NVIDIA H100 GPU and it shows up to 3$\times$ end-to-end speedup while reducing total memory consumption by up to 31% at the decode phase compared to the full KV cache.
@@ 2 Related Work
A number of recent approaches have focused on improving the efficiency of attention mechanisms in LLMs, particularly when dealing with long contexts. One feasible solution is to share KV cache across multiple layers [5] or selectively drop attention for some layers [21]. Others [29, 10] propose mixed attention designs where some layers use global attention while others use local attention. Multi-Query Attention (MQA) [26] and Grouped-Query Attention (GQA) [2] are widely adopted by many recent LLMs [1, 13, 32], which reduce KV cache by sharing them across multiple attention heads. All the above techniques directly modify the attention architecture and require integration since pre-training.
@@
Another direction is to improve attention efficiency with training-free techniques. These techniques can be further categorized as prefill phase acceleration, decode phase acceleration, or accelerating both phases together. For example, StreamingLLM [13] combines initial and local-window attention to reduce the KV cache into a constant size regardless of the sequence length. DuoAttention [35] and RazorAttention [30] improve upon StreamingLLM by applying global attention on retrieval heads and StreamingLLM-style attention on rest heads. While these techniques can be applied to both the prefill and decode phases of LLM inference, other techniques focus on accelerating only one of the two phases.
@@
For prefill phase attention acceleration, MInference [14] identifies three distinct patterns in long-context attention matrices that can be harnessed for efficient sparse operations with customized GPU kernels. SeerAttention [8] explores dynamic block-level sparsity in attention module with a learnable gate.
@@
A common approach for attention acceleration at the decode phase is through permanent KV cache eviction, which can save both KV cache fetching bandwidth and storage requirements. H2O [37] observes that a small subset of tokens, known as heavy-hitters, dominates the attention computation, thus only keeps recent and heavy-hitter tokens. SnapKV [20] employs an observation window at the end of the prompt to identify critical KV tokens of the input prompt. Then it uses a clustering algorithm via pooling to retain critical KV token clusters without loss the completeness of the information. Quest [31] observes that permanent KV cache eviction could lead to inevitable accuracy loss, and proposes query-aware selection on top-k KV tokens based on approximation attention with representative vectors of contiguous key cache pages. On the other hand, SparQ and Loki [25, 28] conduct approximation attention by selecting only important indices on the head dimension instead. The above approaches can save KV cache compute and data fetching from memory but not KV cache storage. InfiniGen [18] tackles this challenge by offloading the entire KV cache to CPU memory and only fetching the selected KV tokens to GPU memory when needed. MagicPIG[6] discovers that using importance sampling is more efficient than top-k estimation and proposes an approximation attention solution leveraging locality sensitive hashing (LSH) and CPU offloading.
@@
In contrast to prior work, RocketKV combines both permanent KV cache eviction and dynamic KV token selection as two consecutive stages for accelerating the decode phase and utilizes the advantages of both worlds. As a result, it achieves remarkable memory bandwidth and storage savings without the need for sophisticated system-level optimizations such as CPU offloading, while maintaining remarkable accuracy even at low token budgets.
@@ 3.1 Observation
As discussed previously in Sections 1 and 2, existing training-free KV cache eviction methods for decode phase acceleration can be categorized into two types: permanent KV cache eviction and dynamic KV token selection. To understand the effectiveness of strategies in each type, Figure 1 presents the accuracy comparison of four different methods for qasper benchmark in LongBench [4] on Mistral-7B-Instruct-v0.2 model, where DuoAttention and SnapKV belong to the first type, and Quest and SparQ belong to the second type. As shown in the figure, the accuracy of all four methods drops significantly as the token budget becomes lower than 1024. In comparison, we present Exact-TopK as an oracle method that performs sparse attention on top-k KV token indices selected based on the exact attention scores. Exact-TopK achieves negligible accuracy drop compared to the Full-KV attention even with a token budget of 256, which indicates that all those practical methods fail to accurately predict top-k KV tokens at low token budgets.
To improve the prediction accuracy, we observe that permanent KV cache eviction and dynamic KV token selection are complementary to each other and can be seamlessly combined together. In fact, we can perform permanent KV cache eviction with a much larger token budget first and then conduct dynamic KV token selection on the remaining KV tokens. By doing this, the chance of permanently evicting important KV tokens is greatly reduced. Meanwhile, dynamic KV token selection only needs to perform on a subset of KV tokens rather than the original full set, significantly lowering the prediction difficulty.
Figure 1: Comparing permanent KV cache eviction and dynamic KV token selection for qasper benchmark.
3.2 RocketKV
Based on the above observation, we propose RocketKV, a two-stage KV cache compression method for decode phase acceleration. As shown in Figure 2, RocketKV performs coarse-grain KV cache eviction at the first stage. The purpose of this stage is to remove KV tokens with low importance while trying to keep the majority of important tokens. In the second stage, it conducts fine-grain dynamic KV token selection on the remaining KV tokens followed by top-k sparse attention. For a given token budget $t$ and associated compression ratio $c$, we evenly split the compression ratio across both stages so that each stage results in a compression ratio of $\sqrt{c}$. To achieve the best performance at each stage, we further propose a highly effective KV cache eviction and dynamic KV token selection method for each stage.
@@ 3.2 RocketKV
3.3 First Stage: SnapKV++
Algorithm 1 SnapKV++
  Input: sequence length $S$, query vector $q$, key tensor $K$, value tensor $V$, kernel sizes $ks_{1}$ and $ks_{2}$, kernel threshold $Thr$, top-k KV indices $k$
  # compute Attention scores based on observation window
  $s_{1}\leftarrow score(q_{obs},K_{pre})$
  # accumulate attention score in seq. and group dim
  $s_{2}\leftarrow sum(s_{1})$
  # adaptively select kernel size for pooling
  $ks\leftarrow ks_{1}$ if $S\geq Thr$ else $ks_{2}$
  # perform pooling in seq. dim
  $s_{3}\leftarrow pool(s_{2},ks)$
  # get indices with top-k attention scores
  $i\leftarrow argtopk(s_{3},k)$
  # generate pruned KV cache
  $K_{cache}\leftarrow K_{[i,obs]},V_{cache}\leftarrow V_{[i,obs]}$
  return $K_{cache},V_{cache}$
Algorithm 2 Hybrid Attention (only contains step 2 and 3)
  Input:query vector $q$, key tensor $K$, value tensor $V$, element-wise max/min key tensor $K_{max}/K_{min}$, top-r $q$ indices $r$, top-k KV indices $k$
  # get top-r indices along head dim from sum of $|q|$ in group dim
  $i_{1}\leftarrow argtopk(sum(|q|,dim=group),r)$
  # get signs of top-r indices from sum of q in group dim
@@ 3.3 First Stage: SnapKV++
  # get indices with top-k attention scores along seq. dim
  $i_{2}\leftarrow argtopk(s_{1},k)$
@@
Figure 3: Illustration of hybrid attention with figure (up) and algorithm (bottom).
In the first stage, we propose SnapKV++ as an efficient KV cache compression method on the input prompt. Compared with the original SnapKV, SnapKV++ introduces two enhancements that greatly improve the performance: 1) Full compatibility to GQA, and 2) adaptive kernel size for pooling.
The original SnapKV method keeps crucial KV tokens based on aggregated per-head attention scores across the observation window. In the case of GQA, each attention head within an attention group will keep a separate set of KV cache tokens, which could introduce redundant storage of the same KV token. To be fully compatible with GQA, SnapKV++ accumulates per-group rather than per-head attention score, and hence the selected KV tokens are shared across the entire attention group.
SnapKV uses pooling along the sequence dimension so that the critical KV tokens are selected along with their neighbor tokens. It demonstrates better accuracy with pooling because it retains the completeness of selected information. The employed kernel sizes for pooling are quite small (e.g., a kernel size of 7 for LongBench [4]). Since SnapKV++ is only used for coarse-grain KV token eviction at the first stage, we discovered that the optimal kernel sizes for pooling are much larger. In fact, the optimal kernel size can grow as the input sequence length increases. Therefore, SnapKV++ provides a simple adaptive kernel size selection mechanism based on the input sequence length: it selects a smaller kernel size if the input sequence length is shorter than a threshold value, otherwise, it will switch to a larger kernel size. Those two kernel sizes and threshold values are empirically determined in our work as shown in Section 4.4. We found this mechanism is sufficient to capture close-to-optimal performance in our studies, so we did not explore more sophisticated algorithms.
The overall algorithm of SnapKV++ is illustrated in Algorithm 1. We found that with those two enhancements, SnapKV++ greatly outperforms the original SnapKV method as shown in Section 4.4.
3.4 Second Stage: Hybrid Attention

### v2

@@ Abstract
Transformer-based Large Language Models rely critically on the KV cache to efficiently handle extended contexts during the decode phase. Yet, the size of the KV cache grows proportionally with the input length, burdening both memory bandwidth and capacity as decoding progresses. To address this challenge, we present RocketKV, a training-free KV cache compression strategy containing two consecutive stages. In the first stage, it performs coarse-grain permanent KV cache eviction on the input sequence tokens. In the second stage, it adopts a hybrid sparse attention method to conduct fine-grain top-k sparse attention, approximating the attention scores by leveraging both head and sequence dimensionality reductions. We show that RocketKV provides a compression ratio of up to 400$\times$, end-to-end speedup of up to 3.7$\times$ as well as peak memory reduction of up to 32.6% in the decode phase on an NVIDIA A100 GPU compared to the full KV cache baseline, while achieving negligible accuracy loss on a variety of long-context tasks. We also propose a variant of RocketKV for multi-turn scenarios, which consistently outperforms other existing methods and achieves accuracy nearly on par with an oracle top-k attention scheme.
Machine Learning, ICML, LLM, KV cache, Compression
@@ 1 Introduction
In Transformer-based LLM inference (Vaswani et al., 2017; Pawar et al., 2024), the key-value cache (KV cache)—which stores past attention keys and values to avoid recomputation—becomes a major bottleneck during the decode phase, as its size scales linearly with both the sequence length and batch size. For example, the Llama3.1-70B-Instruct (AI@Meta, 2024) model with a batch size of 32, and a context length of 32K requires around 320GB of KV cache storage at FP16 precision, which even advanced hardware (NVIDIA, 2024; Jouppi et al., 2023; AMD, 2024) can hardly handle.
@@
Fortunately, previous work has shown that only a small subset of KV tokens is required at each decode step to maintain accuracy (Zhang et al., 2023; Tang et al., 2024; Ribar et al., 2024; Li et al., 2024; Singhania et al., 2024; Ge et al., 2024). Therefore, if those KV tokens can be accurately predicted in advance, dense attention operations can be replaced with sparse attention operations with significant memory bandwidth and capacity improvement. These methods often fall into two categories: 1) permanent KV token eviction, and 2) dynamic KV token selection. The former results in both memory bandwidth and storage savings, but could lead to noticeable accuracy loss if KV tokens dropped earlier are needed by later decode steps. The latter avoids this shortcoming by keeping all KV tokens in memory and dynamically selecting a subset each time. Hence, it only results in memory bandwidth savings, but often requires extra memory storage overhead for auxiliary data.
Figure 1: Existing KV token dropping methods fail to match the accuracy scores of oracle top-k attention (Exact-TopK) on Mistral-7B-Ins-v0.2 in the qasper benchmark.
To understand the effectiveness of existing KV token dropping methods, Figure 1 presents the accuracy comparison of four different methods (two in each category) for qasper benchmark in LongBench (Bai et al., 2024) on the Mistral-7B-Instruct-v0.2 model (Jiang et al., 2023). We observe that the accuracy of all four methods drops significantly as the token budget becomes lower than 1024 while an oracle top-k attention scheme achieves negligible accuracy drop even with a token budget of 256. This indicates that existing practical methods fail to accurately predict top-k KV tokens under low token budgets.
@@
To improve the prediction accuracy, we propose RocketKV, a two-stage KV cache compression method that combines permanent KV token eviction with dynamic KV token selection to accelerate the decode phase of LLM inference. The RocketKV framework enables flexible integration of a wide range of KV cache compression techniques at each stage. To achieve optimal performance, we directly adopt an existing method SnapKV (Li et al., 2024) for coarse-grain permanent KV token eviction in the first stage. In the second stage, we propose hybrid sparse attention (HSA) to perform fine-grain dynamic KV token selection, which estimates KV token indices with top-k attention scores via two-dimensional reductions. To balance the compression between two stages, RocketKV introduces an adaptive compression decomposition mechanism that intelligently splits any given target compression ratio across two stages. Combining these two stages together, RocketKV achieves significant memory bandwidth and storage savings with negligible accuracy loss across a wide variety of models and downstream tasks.
Moreover, recent work (Li et al., 2025) demonstrates that permanent KV token eviction suffers in multi-turn decoding because important KV tokens vary significantly across queries. To address this challenge, we propose a variant of RocketKV called RocketKV-MT for multi-turn scenarios where we do not evict the unselected KV tokens in the first stage but keep them all for later turns. Yet, the decode phase is still restricted to perform dynamic selection on the filtered KV tokens from the first stage in each turn. By doing this, RocketKV-MT achieves the same memory traffic savings as RocketKV but does not introduce memory storage savings.
In summary, we make the following contributions:
@@
  • We analyze and identify the limitations of existing KV token dropping methods and then propose RocketKV as a two-stage KV cache compression scheme. We further introduce a variant RocketKV-MT for multi-turn decoding scenarios.
  • We propose a hybrid sparse attention (HSA) for dynamic KV token selection with two-dimensional reductions in the second stage while directly adopting SnapKV in the first stage. We design an adaptive compression decomposition mechanism to balance the compression between the two stages.
@@
  • We conduct a comprehensive evaluation of RocketKV and RocketKV-MT on a wide variety of models and downstream tasks. RocketKV consistently demonstrates comparable accuracy with full KV attention at up to 400$\times$ compression ratio, while achieving up to 3.7$\times$ end-to-end speedup and 32.6% peak memory saving at the decode phase on an NVIDIA A100 GPU. Meanwhile, RocketKV-MT is more suitable for multi-turn decoding and performs on par with oracle top-k attention.
@@ 2 Related Work
A number of recent approaches have focused on improving the efficiency of attention mechanisms in LLMs, particularly when dealing with long contexts. One feasible solution is KV cache sharing across multiple layers (Brandon et al., 2024) or selectively dropping attention for some layers (Ling et al., 2024). Others (Sun et al., 2024; Ho et al., 2024) propose mixed attention designs where some layers use global attention while others use local attention. Multi-Query Attention (MQA) (Shazeer, 2019), Grouped-Query Attention (GQA) (Ainslie et al., 2023), and Multi-head Latent Attention(MLA) (Liu et al., 2024a)are widely adopted by many recent LLMs (AI@Meta, 2024; Jiang et al., 2023; Team et al., 2024; Liu et al., 2024b), which reduce the KV cache by sharing key-value pairs across multiple attention heads. All above techniques directly modify the attention architecture and require integration since pre-training.
@@
Another direction is to improve attention efficiency with training-free techniques. These techniques can be further categorized as prefill phase acceleration, decode phase acceleration, or accelerating both phases together. For example, StreamingLLM (Jiang et al., 2023) combines initial and local-window attention to reduce the KV cache into a constant size regardless of the sequence length. DuoAttention (Xiao et al., 2025) and RazorAttention (Tang et al., 2025) improve upon StreamingLLM by applying global attention on retrieval heads and StreamingLLM-style attention on rest heads. While these techniques can be applied to both the prefill and decode phases, other techniques focus on accelerating only one of the two phases.
@@
For prefill phase acceleration, MInference (Jiang et al., 2024) identifies three distinct patterns in the attention matrix that can be harnessed for efficient sparse operations with customized GPU kernels. SeerAttention (Gao et al., 2024) explores dynamic block-level sparsity in the attention module with a learnable gate. Meanwhile, XAttention (Xu et al., 2025) leverages the sum of antidiagonal values in the attention matrix for estimating block importance.
@@
A common approach for attention acceleration at the decode phase is through permanent KV cache eviction, which can save both memory bandwidth and storage requirements. H2O (Zhang et al., 2023) observes that a small subset of tokens, known as heavy-hitters, dominates the attention computation, thus only keeps recent and heavy-hitter tokens. SnapKV (Li et al., 2024) employs an observation window at the end of the prompt to identify critical KV tokens of the input prompt. Then it uses a clustering algorithm via pooling to retain critical KV token clusters without losing information completeness. Ada-KV (Feng et al., 2024) proposes an adaptive budget allocation strategy to provide better token budget utilization across individual heads.
Meanwhile, Quest (Tang et al., 2024) observes that permanent KV cache eviction could lead to inevitable accuracy loss and proposes query-aware selection on top-k KV tokens based on approximation attention with representative vectors of contiguous key cache pages. On the other hand, SparQ and Loki (Ribar et al., 2024; Singhania et al., 2024) conduct approximation attention by selecting only important indices on the head dimension instead. The above approaches can save KV cache compute and data fetching from memory, but not KV cache storage. InfiniGen (Lee et al., 2024) tackles this challenge by offloading the entire KV cache to CPU memory and only fetching the selected KV tokens to GPU memory when needed. MagicPIG(Chen et al., 2025) discovers that using importance sampling is more efficient than top-k estimation and proposes an approximation attention solution leveraging locality sensitive hashing (LSH) and CPU offloading.
@@
In contrast to prior work, RocketKV combines both permanent KV cache eviction and dynamic KV token selection as two consecutive stages for accelerating the decode phase and utilizes the advantages of both worlds. As a result, it achieves remarkable memory bandwidth and storage savings without the need for sophisticated system-level optimizations such as CPU offloading.
@@ 3 Proposed Method: RocketKV
Figure 2: The CDF of maximum sequence length and number of unique top-k indices (k=256) across all decoding steps for 200 questions in the qasper benchmark. Data was collected from a random head (layer 31 head 0) in Mistral-7B-Instruct-v0.2.
@@ 3.1 Observation
As shown previously in Figure 1, existing token dropping methods, regardless of permanent KV token eviction or dynamic KV token selection, fail to match the accuracy of oracle top-k attention (Exact-TopK) under low token budgets. To further understand what causes the accuracy mismatch, we analyze a random attention head (layer 31 head 0) from Mistral-7B-Instruct-v0.2 and present the cumulative distribution function (CDF) of both the maximum sequence length and number of unique KV indices selected by Exact-TopK (k=256) across all decoding steps in the qasper benchmark. The reason behind this analysis is that in order to match the accuracy of Exact-TopK, we need to keep all important KV tokens that are selected by at least one top-k attention operation across all decoding steps. As shown in Figure 2, although the maximum sequence length can reach as high as 25000, the number of unique top-k indices only goes up to 1200. This implies that, ideally, a permanent KV token eviction method should be able to close the accuracy gap under a token budget of 1200. To further reduce the token budget, we realize that dynamic KV token selection can be applied on the filtered KV token set after permanent KV token eviction. Since the filtered set is much smaller than the original full KV cache, the difficulty for accurate top-k prediction is greatly reduced. Therefore, an ideal solution would be to perform permanent KV cache eviction with a larger token budget first and then conduct dynamic KV token selection on the remaining KV tokens. This fusion evicts unimportant tokens and also makes the dynamic selection more accurate, motivating our RocketKV design.
@@
3.2 RocketKV Overview
Based on our observation, we propose RocketKV, a two-stage KV cache compression method for decode phase acceleration. As shown in Figure 3, RocketKV performs coarse-grain KV cache eviction in the first stage. The purpose of this stage is to remove KV tokens with low importance while keeping the majority of important tokens. In the second stage, it conducts fine-grain dynamic KV token selection on the remaining KV tokens, followed by top-k sparse attention. The RocketKV framework is generic, and many existing KV cache compression methods can fit into the corresponding stage. For example, SnapKV (Li et al., 2024) or Ada-KV (Feng et al., 2024) can be used for the first stage, and Quest (Tang et al., 2024) or SparQ (Ribar et al., 2024) can be applied to the second stage. To achieve the best performance at each stage, we directly adopt SnapKV in the first stage while proposing a hybrid sparse attention (HSA) method in the second stage.
3.3 First Stage: SnapKV
In the first stage, we directly adopt SnapKV for permanent KV token eviction on the input KV tokens. SnapKV’s key idea is to rely on the aggregated attention scores between the input context and observation window in the end to select the most relevant tokens to keep within the input prompt. The original SnapKV method selects crucial KV tokens on a per attention head basis. In case of grouped-query attention (GQA), each attention head within an attention group keeps a separate set of KV cache tokens, which could introduce redundant storage of the same KV token. To reduce KV token storage with GQA, we follow the Ada-KV work to perform token selection on a per-group basis according to aggregated per-group attention scores so that the selected KV tokens are shared across the entire attention group. SnapKV uses pooling along the sequence dimension to ensure critical KV tokens are selected along with their neighbor tokens. It demonstrates better accuracy with pooling because it retains the completeness of selected information. The employed kernel sizes for pooling are quite small (e.g., a kernel size of 7 for LongBench (Bai et al., 2024)). Since in our case, SnapKV is only used for coarse-grain KV token eviction at the first stage, we discovered that the optimal kernel sizes for pooling are much larger. We empirically set the kernel size to 63 in all our experiments.
Algorithm 1 HSA Algorithm (only contains step 2 and 3)
  Input:query vector $q$, key tensor $K$, value tensor $V$, element-wise max/min key tensor $K_{max}/K_{min}$
  # get top-$k_{1}$ indices along head dim from sum of $|q|$ in group dim
  $i_{1}\leftarrow argtopk(sum(|q|,dim=group),k_{1})$
  # get signs of top-$k_{1}$ indices from sum of q in group dim
@@ 3.3 First Stage: SnapKV
  # get indices with top-$k_{2}$ attention scores along seq. dim
  $i_{2}\leftarrow argtopk(s_{1},k_{2})$
@@
Figure 4: Illustration of hybrid sparse attention with figure (left) and algorithm (right).
3.4 Second Stage: Hybrid Sparse Attention